OpenStudy (anonymous):

Say that you (or your parents) are purchasing a used car for $19,850. The sales tax is 7.5%, the down payment is $1,000.00, and you have an average credit rating. If your first payment is $425.98, how much of the payment goes toward the principal?

OpenStudy (anonymous):

add the sales tax then take off the down payment

OpenStudy (anonymous):

19850*.075=1488.75 19850+1488.75=21338.75 21338.75-1000=20338.75

OpenStudy (anonymous):

you are making payments on the 20338.75

OpenStudy (anonymous):

that is not the answer though. that is just getting you started @therussiancircus

OpenStudy (anonymous):

you can set it up like this \[\frac{ 425.98 }{ 20338.75 }=\frac{ x }{ 19850 }\]

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

that will be how much is going towards the principle

OpenStudy (anonymous):

and ill give you a hint it is between 400 and 425

OpenStudy (anonymous):

Do you cross multiply?